OpenStudy (anonymous):

Write a trinomial of degree 4 such that the GCF of its terms is 1. Whaat??

OpenStudy (amistre64):

its prime, it cant be factored

OpenStudy (amistre64):

x^4 + 1 seems to fit just fine

OpenStudy (amistre64):

trinomial .. 3 terms tho

OpenStudy (anonymous):

Idk what you're saying..

OpenStudy (amistre64):

find a trinomial of degree 4 that cant be factored - has no real roots

OpenStudy (anonymous):

I dont know what that means... what is a trinomial

OpenStudy (amistre64):

(1+i-x)(1-i-x)(2+i-x)(2-i-x) should produce one

OpenStudy (amistre64):

a trinomial is an equation that only has 3 term in it

OpenStudy (amistre64):

ugh, that last one has 5 terms

OpenStudy (anonymous):

im so lost

OpenStudy (amistre64):

consider the roots: a+-bi and c+-di such that:\[f(x)=(a+bi-x)(a-bi-x)(c+di-x)(c-di-x)=x^4+px^3+qx+k\] such that either p or q is zero

OpenStudy (anonymous):

....Whhhaaaaaaaaaaaaaaaaaaaaaaaat????????????

OpenStudy (amistre64):

\[a^2 b^2-2 a^2 b x+a^2 x^2+a^2-2 a b^2 x+4 a b x^2\\~~~~-2 a x^3-2 a x+b^2 x^2+b^2-2 b x^3-2 b x+x^4+2 x^2+1\] and gather like terms ... to compare coefficients with

OpenStudy (anonymous):

trinomial -> means 3 terms. degree 4 means the highest power ( we will take x^4) for example: x^4 + x^3 + x^2

OpenStudy (anonymous):

... umm, that seems.. like not what im supposed to do

OpenStudy (anonymous):

im so confused.

OpenStudy (anonymous):

are the coefficients GCF is indeed 1 ?

OpenStudy (anonymous):

Your questions was not gramatically correct, so i have no idea what you just said

OpenStudy (amistre64):

GCF of terms, not of coefficients

OpenStudy (anonymous):

would it just be 4^3?

OpenStudy (anonymous):

i believe they meant GCF of coefficients

OpenStudy (amistre64):

asking for answers like that can lead to suspensions

OpenStudy (anonymous):

Thats not what i mean..

OpenStudy (amistre64):

i cant seem to develop a simple solution from my idea